Multiply 53 by 24, firstly by breaking down 24 in its terms (20 + 4) and secondly by breaking down 24 in its factors (6 * 8). Show all your steps.


1
Expert's answer
2021-08-23T16:16:59-0400
"53\times24=53\times(20+4)=53\times20+53\times4"

"=1060+212=1272"

"53\times24=53\times(3\times8)=(53\times3)\times8"

"=159\times8=1272"